OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, 2) with slope 3.

OpenStudy (anonymous):

sub in what you know. y=mx+b where m=3, x=4, and y=2. so 2=3(4)+b 2=12+b 2-12=b b=-10. now you have your m and b, the 2 things you need for your equation. so y=mx+b y=2x-10.

OpenStudy (anonymous):

Thanks a lot~! that helps me out more than you know :)